# Decision-aware training of spatiotemporal forecasting models to select a top K subset of sites for intervention

### v2

@@ Appendix A Pseudocode for Training
Algorithm A.1 Decision-aware ML training for top-K tasks judged by BPR
@@ Appendix B Ranking Demo for BPR
BPR fraction of trials each site in top $K{=}3$ of $\bm{r}$
$K{=}1$ $K{=}3$ $K{=}6$ A:#1 A:#2 A:#3 B:#4 B:#5 B:#6 C:#7 C:#8 C:#9
@@ Appendix C Experimental Details and Results from Real-World Data
Aransas TX Whooping Cranes 1338 boxes, each 500m×$\times$500m 60 years, 1952-2011 (last 10 years used) Count of bird spotted in aerial survey Past observations, Location, Time, Month
@@
C.1 Opioid-related Overdose Forecasting Results
C.1.1 Features
@@
  • Overdose gravity, an weighted average of the prior year’s overdose deaths in all contiguous tracts. We construct the feature in the same way as (Marks et al., 2021a). Note that the original paper describes a weighted average over all regions within a radius, but the provided code uses only immediately contiguous locations. We follow the implementation from the code.
  • Covariates from the Social Vulnerability Index (SVI) (CDC ATSDR, 2018). These covariates are available as 5-year estimates for every census tract in the United States and are updated every 2 years. The SVI measures report the percentile ranking of every census tract according to 4 themes: Socioeconomic, Household Composition & Disability, Minority Status & Language, and Housing Type & Transportation. We use 5 variables: each tract’s ranking in each of the four themes as well as its composite ranking.
@@ C.1.2 Hyperparameter ranges
  • Number of samples for SF estimator: 100
  • Number of samples for Perturb estimator: 100
@@
C.2 Endangered Bird Forecasting Results
C.2.1 Features
@@
C.2.2 Hyperparameter ranges
The same hyperparameter and model selection was performed on Whooping Crane data as the opioid experiment in Section  C.1.
@@ Appendix E Model for Negative Binomial Mixed-Effects
Here, $\sigma_{0}$ is the standard deviation of the random intercepts, $\sigma_{1}$ is the standard deviation of the random slopes, and $\rho$ is the correlation between the random intercepts and slopes, where $-1\leq\rho\leq 1$. We pack these hyperparameters into a separate vector $\eta=\{\sigma_{0},\sigma_{1},\rho\}$.

### v3

@@ Appendix A Pseudocode for Training
Algorithm A.1 Decision-aware ML training
@@ Appendix B Ranking Demo for BPR
B.1 Justification for ratio estimator
In the text, we propose the ratio estimator for ranking locations: $\bm{r}=\mathbb{E}[\frac{\bm{y}}{\bm{1}\cdot\bm{y}}]$ and justify its use by demonstrating that it is a better bound on our decision loss BPR than the mean estimator. However, a natural question is, why not use a ranking estimator which more closely resembles BPR, such as $\bm{r}=\mathbb{E}[\frac{\bm{y}}{\text{TopKMask}(\bm{y},K)}\cdot\bm{y}]$. In practice, this estimator is equivalent to the ratio estimator in expectation, and will select the same top-K locations. Accordingly, we use the ratio estimator for its improved computational speed and simplicity.
B.2 Demo experiment
@@
BPR fraction of trials each site in top K=3K{=}3 of $\bm{r}$
K=1K{=}1 K=3K{=}3 K=6K{=}6 A:#1 A:#2 A:#3 B:#4 B:#5 B:#6 C:#7 C:#8 C:#9
@@ Appendix C Experimental Details and Results from Real-World Data
Aransas TX Whooping Cranes 1338 boxes, each $\times500m$ 60 years, 1952-2011 (last 10 years used) Count of bird spotted in aerial survey Past observations, Location, Time, Month
@@
C.1 Results on alternative models and metrics
ANWR TX Cranes MAE RMSE BPR-50
Last timestep 0.24 1.04 0.27
Avg over 10 0.25 0.78 0.38
Chance decision, $\hat{y}=0$ 0.15 0.79 0.05
EpiGNN 0.38 0.70 0.06
PG 1.03 4.85 0.35
SPO+ 0.15 0.83 0.18
NLL Only 0.35 1.14 0.38
DAML ($\epsilon=1$) 9.08 35.89 0.39
BPR Only 40495 40495 0.41
(a) ANWR TX Cranes 2009-2010.
Cook County IL Overdose MAE RMSE BPR-100
Last timestep 1.07 1.66 0.76
Avg over 5 0.99 1.58 0.80
Chance decision, $\hat{y}=0$ 1.37 2.46 0.20
EpiGNN 1.33 2.03 0.32
PG 5.65 17.03 0.80
SPO+ 1.25 2.38 0.74
NLL Only 1.03 1.58 0.78
DAML ($\epsilon=1$) 1.12 1.84 0.80
BPR Only 70.31 152.72 0.82
(b) Cook County IL 2021-2022.
Table C.2: Performance comparison between different model families for error-based and decision metrics. The top 2 rows represent simple historical baselines: using either the previous timestep alone, or an average over a larger amount (10 bi-months for the crane dataset, and 5 years for Cook County). Next, the Chance decision model chooses spatial locations by random chance, and uses all 0’s for predicted $\hat{y}$ in MAE and RMSE calculations. Despite the simplicity, this is a competitive model on the error-based metrics due to sparsity. It outperforms all others on MAE on the Cranes dataset. Next is EpiGNN, a spatiotemporal forecasting model based on graph neural networks, trained to minimize RMSE. Although this model has the best performing RMSE on the crane dataset, it makes poor top-K decisions on both. Finally, the last three rows show different ways to train the negative binomial mixed effects regression model in 14. NLL Only trains to maximize likelihood alone (8), BPR Only is our direct-loss minimization trained to maximize BPR alone (9), and DAML (13) is our hybrid loss that allows a user to explore the pareto frontier between likelihood and BPR-K.
C.2 Opioid-related Overdose Forecasting Results
C.2.1 Features
@@
  • Overdose gravity, an weighted average of the prior year’s overdose deaths in all contiguous tracts. We construct the feature in the same way as (marks_identifying_2021). Note that the original paper describes a weighted average over all regions within a radius, but the provided code uses only immediately contiguous locations. We follow the implementation from the code.
  • Covariates from the Social Vulnerability Index (SVI) (cdcatsdrCDCATSDRSocial2022). These covariates are available as 5-year estimates for every census tract in the United States and are updated every 2 years. The SVI measures report the percentile ranking of every census tract according to 4 themes: Socioeconomic, Household Composition & Disability, Minority Status & Language, and Housing Type & Transportation. We use 5 variables: each tract’s ranking in each of the four themes as well as its composite ranking.
@@ C.2.2 Hyperparameter ranges
  • Number of samples for score function trick estimator: 100
  • Number of samples for perturbation estimator: 100
@@
C.3 Endangered Bird Forecasting Results
C.3.1 Features
@@
C.3.2 Hyperparameter ranges
The same hyperparameter and model selection was performed on Whooping Crane data as the opioid experiment in Section  C.2.
@@ Appendix E Model for Negative Binomial Mixed-Effects
Here, $\sigma_{0}$ is the standard deviation of the random intercepts, $\sigma_{1}$ is the standard deviation of the random slopes, and $\rho$ is the correlation between the random intercepts and slopes, where $1\leq\rho\leq 1$. We pack these hyperparameters into a separate vector $\eta=\{\sigma_{0},\sigma_{1},\rho\}$.